Rico is twice old as Ryan.In four years,Rico's age will be three times as Ryan's age is six years ago.How old are they now?
Answer by Maths68(1474)

Let
Present age of Rico = r
Present age of Ryan = y


Rico is twice old as Ryan
r=2y........(1)
In four years,Rico's age will be three times as Ryan's age is six years ago
Rico age in four years = r+4
Ryan age six years ago = y-6
r+4=3(y-6)
r+4=3y-18
r=3y-18-4
r=3y-22.............(2)
Put the value of r in (2) from (1)
2y=3y-22
2y-3y=-22
-y=-22
Multiply by -1 both sides
y=22
Put the value of y in (1)
r=2y
r=2(22)
r=44
Present age of Rico = r = 44 years old
Present age of Ryan = y = 22 years old
Six years ago Ryan was = 22-6 = 16
In four years Rico will be = 44+4 = 48 (which is 3 times of 16)